# Multi-Agent Trajectory Prediction with Difficulty-Guided Feature Enhancement Network

### v2

@@ Abstract
Trajectory prediction is crucial for autonomous driving as it aims to forecast the future movements of traffic participants. Traditional methods usually perform holistic inference on the trajectories of agents, neglecting the differences in prediction difficulty among agents. This paper proposes a novel Difficulty-Guided Feature Enhancement Network (DGFNet), which leverages the prediction difficulty differences among agents for multi-agent trajectory prediction. Firstly, we employ spatio-temporal feature encoding and interaction to capture rich spatio-temporal features. Secondly, a difficulty-guided decoder is used to control the flow of future trajectories into subsequent modules, obtaining reliable future trajectories. Then, feature interaction and fusion are performed through the future feature interaction module. Finally, the fused agent features are fed into the final predictor to generate the predicted trajectory distributions for multiple participants. Experimental results demonstrate that our DGFNet achieves state-of-the-art performance on the Argoverse 1&2 motion forecasting benchmarks. Ablation studies further validate the effectiveness of each module. Moreover, compared with SOTA methods, our method balances trajectory prediction accuracy and real-time inference speed. The code is available at https://github.com/XinGP/DGFNet.
@@ I Introduction
Trajectory prediction is a crucial component of current autonomous driving systems. Its goal is to infer the future trajectory distribution of agents based on historical information within the scene. In addition to the historical motion information of the agents, various complex factors must be considered, such as lane markings and traffic light constraints in the map, as well as social interactions between agents. These considerations ensure that the future trajectory distribution can accurately reflect the agents’ movement trends, thereby ensuring the proper functioning of the autonomous driving system.
@@
Currently, mainstream prediction model architectures can be roughly divided into three steps: feature encoding, feature interaction, and feature decoding. Additionally, some methods improve the accuracy and robustness of prediction models through feature enhancement [1, 2, 3, 4, 5, 6, 7, 8, 9]. These methods involve intermediate modeling of future intentions or trajectories and interact and fuse these with the original encoded features. Although these methods significantly improve model performance compared to mainstream model architectures, they neglect the inherent heterogeneity in prediction difficulty among agents, i.e., the natural differences in prediction difficulty for different agents within the scene. As illustrated in the Fig. 1, The primary difference in our pipeline is the addition of an intermediate module that generates reliable future trajectories, thereby better prediction results can be achieved for agents with higher prediction difficulty.
In real driving scenarios, human drivers often predict the future behaviors of most agents subconsciously, relying on intuition. However, when faced with agents that may pose risks and exhibit high prediction difficulty, human drivers tend to carefully consider their behavior and motivations [10]. Inspired by human drivers, we aim to investigate the effectiveness of prediction methods that follow the principle of easy-to-hard in multi-agent trajectory prediction tasks.
The main contributions of this paper can be summarised as follows:
@@
II Related Work
Future Feature Enhancement. Compared to conventional pipelines, methods that enhance features using future information allow networks to incorporate future constraints and interactions. These methods can be broadly categorized into two approaches: intention and future trajectories. TNT [1] and PECNet [2] model vehicle intention as latent variables and generate different prediction trajectories conditioned on vehicle intentions, forming predicted trajectories under certain constraint conditions. GANet [3] and ADAPT [4] integrate and fuse vehicle intentions for future feature enhancement through feature interaction. Similarly, Prophnet [5] and FFINet [6] model intermediate latent variables in the form of trajectories, where the output of intermediate decoders is one or multiple trajectories. It should be noted that the first prediction results of both Prophnet and FFINet are obtained by decoding only the historical trajectory features, which leads to a large error in both first prediction results. Additionally, there is a method that models future trajectories through planner sampling, followed by feature enhancement [8, 9].
Most relevant to our work is M2I [7], which exploits potential game relationships between vehicles. It uses a Relation Predictor and a Marginal Trajectory Predictor to input the future trajectories of the leader in the game relationship as prior information into the Conditional Trajectory Predictor. However, we find that this method is highly dependent on the accuracy of the inferred game relationships and future trajectories. Therefore, we use the more easily extractable prediction difficulty as a basis for personalized classification.
Scene Representation. Broadly, vectorization can be categorized into two types: scene-centric and agent-centric. Since the pioneer of vectorization, Vectornet, was designed for single-agent trajectory prediction tasks, it naturally took the last frame coordinates of the vehicle to be predicted as the scene center, and expressed the trajectories and map vector coordinates of all other agents around the scene center. The benefit of this approach is that it enables the preservation of global relative pose features of various objects in the scene during feature extraction, and facilitates the embedding of some prior information or trajectory anchors [5, 1].
However, due to the demands of multi-agent trajectory prediction tasks, there has been a shift towards agent-centric scene representation methods. Building upon agent-centric approaches, some methods aim to preserve as much information as possible about the mutual spatial relationships of various objects in the scene [11, 12, 13, 14, 15]. To achieve this, they incorporate local pairwise-relative pose features and extract and fuse the relative pose features through networks. However, such methods suffer from the limitation that historical and predicted trajectories only retain local motion characteristics, making it inconvenient to embed anchor and prior information. Therefore, we aim to preserve the advantages of each method by integrating the two approaches.
III Method
III-A Overview
@@
(1)Spatio-temporal Feature Extraction. This component extracts spatio-temporal features separately from agent-centric historical trajectory information and scene-centric information using both agent-centric and scene-centric approaches. (2)Feature Interaction. The difference with spatio-temporal feature extraction is that the feature extraction component is compatible for both scenarios. However, feature interaction is not. The feature interaction component is categorized into two types: multi-head attention module and multi-head attention module with edge features. The two interaction methods are respectively applicable to scene-centric and agent-centric scenario representations. (3)Trajectory Decoder. This component is divided into a difficulty-guided decoder and a final decoder. The difficulty-guided decoder operates during the intermediate process, while the final decoder outputs the final prediction results.
@@
Figure 3: Our spatio-temporal feature extraction includes two sets of independent encoders, which extract features based on scene representations (bottom). Subsequently, the extracted features pass through their respective feature interaction modules to obtain interacted Actor features. Following this, we perform future feature enhancement by interacting reliable future trajectories with Actor and Map features. Finally, these inputs are fed into the final predictor to obtain predicted trajectories and their corresponding probabilities.
III-B Problem formulation
Building on the majority of previous work, such as [16, 17, 1, 13, 18], we assume that upstream perceptual tasks can provide high-quality 2D trajectory tracking data for agents in a coordinate system. At the same time, localisation and mapping tasks can provide accurate self-vehicle positioning data and comprehensive HD map data for the urban environment. In other words, for $j$ agents within a given scene, we obtain $x$ and $y$ positions, denoted as $\mathcal{X}^{0:j}_{-t_{h}:0}$, corresponding to the time stamp horizon $\left\{-t_{h},...,0,1,...,t_{f}\right\}$. The trajectory prediction task consists of predicting the future trajectories $\mathcal{Y}^{0:j}_{1:t_{f}}$ of the agents by using the HD map information $\mathcal{M}$ (including road coordinates, topological links, traffic lights ) within the given scene and the historical trajectories output by the tracking task.
We represent the trajectory information of agents and lane details in vectorized form. To elaborate further, for a given agent $i$ within the scenario, its trajectory information, denoted as $\mathcal{X}^{i}$, is represented as a matrix $\mathcal{X}^{i}_{-t_{h}:0}$ comprising a spatio-temporal sequence$\left\{s^{i}_{-t_{h}},s^{i}_{-t_{h}+1},...,s^{i}_{0}\right\}$ over the past $t_{h}$ time step. Similarly, map information is also segmented into predefined sequence vectors to capture various scene details. Analogous to trajectory vectors, map information is represented as $\mathcal{M}_{1:N}=\left\{s_{1},s_{2},...,s_{N}\right\}$, where $N$ denotes the total vector length. Within each lane vector $s_{i}$, there is an inclusion of lane slice coordinates, lane type (e.g., straight or left-turn lane), and map lane details such as traffic signals. This encapsulates diverse map information pertaining to each lane.
@@ III-B Problem formulation
III-C Spatio-temporal feature extraction
In the problem formulation, we explain the process of representing agent and high-definition (HD) map data as vectors, establishing a corresponding mapping between continuous trajectories, map annotations, and sets of vectors. This vectorization method allows us to encode trajectory and map information as vectors. Following the trajectory feature extraction method of LaneGCN, our trajectory feature extraction module mainly uses 1D CNN and downsampling techniques to encode the historical trajectories of all vehicles in the scene, thus obtaining encoded historical trajectory features. For map tensor, we use the PointNet [19] to encode lane nodes and structural information to obtain map features. It is worth mentioning that when processing scene-centric vector information, we do not mix vehicle historical trajectories and lane line information. Our view is that vehicle trajectories, as dynamic vectors, need to be distinguished from static maps to better capture the local motion characteristics of the vehicles. For the extraction of historical trajectory features are described by this:
$\mathcal{Z}_{i}=\mathbf{Conv1d}(\mathbf{Res1d}(\mathcal{X}_{i}))\quad\text{for }i=0,1,\dots,n_{\text{fpn}}-1$ (1)
$\hat{\mathcal{X}}_{i}=\mathbf{Inter}(\mathcal{Z}_{i+1},\text{scale\_factor}=2)+\mathcal{Z}_{i}$ (2)
after Res1d and Conv1d encoding, feature alignment is performed by scaling and interpolation operations. For the agent-centric trajectory tensor, the same feature extraction is used to obtain $\hat{\mathcal{A}}$.
For map feature extraction is described by this:
${\mathcal{H}}_{1}$ $=\mathbf{PAB}_{1}\left(\mathbf{ReLU}(\mathbf{LayerNorm}(\mathcal{M}{W}_{p}+{b}_{p}))\right)$ (3)
$\hat{\mathcal{M}}$ $=\mathbf{PAB}_{2}({\mathcal{H}}_{1})$ (4)
where $\mathcal{M}\in\mathbb{R}^{N_{\text{lane}}\times 10\times C_{\text{in}}}$ are the input lane features, ${W}_{p}\in\mathbb{R}^{C_{\text{in}}\times h}$ and $\mathbf{b}_{p}\in\mathbb{R}^{h}$ are the weights and biases for the projection layer, and PAB represents the PointAggregateBlock.
Similarly, we can obtain the agent-centric map feature $\hat{\mathcal{L}}$. For $\mathcal{RPE}$ we use MLPs for feature extraction to get $\hat{\mathcal{RPE}}$.
III-D Feature interaction
Our intention was to utilize the global representation of scence-centric scenes to embed future trajectory features. Interestingly we found that even without any future enhancement, just concatenating the Actor features after the two interactions gives a better boost. Specific results can be seen later in the ablation experiments. Here we will present the two interaction methods: multi-head attention module and multi-head attention module with edge features, respectively.
@@ III-D Feature interaction
$\mathbf{MHA}(Q,K,V)=\mathbf{softmax}\left(\frac{QK^{T}}{\sqrt{dim_{k}}}\right)V$ (5)
$Q,K,V=W^{q},W^{k},W^{v}$ (6)
where $\mathbf{MHA}$ denotes the Multi-Head Attention operation. $Q$, $K$, $V$ are computed by linear projections $W^{q}$, $W^{k}$, $W^{v}$ applied to input vectors. The attention mechanism computes scaled dot-product attention using $Q$ and $K$, and applies it to $V$ after softmax normalization.
The feature interaction module for carrying edge features is described as follows:
@@
$\mathbf{MHA}(Q^{\prime},K^{\prime},V^{\prime})=\mathbf{softmax}\left(\frac{Q^{\prime}K^{\prime T}}{\sqrt{dim_{k}}}\right)V^{\prime}$ (7)
$Q^{\prime},K^{\prime},V^{\prime}=Q+W^{q}\text{edge}^{\prime},K+W^{k}\text{edge}^{\prime},V+W^{v}\text{edge}^{\prime}$ (8)
@@
here, $Q^{\prime},K^{\prime},V^{\prime}$ are the query, key, and value tensors after linear transformations that incorporate edge features, and $\text{edge}^{\prime}$ is the edge feature after linear transformation.
Finally, the normalized output after attention and dropout application is:
$\hat{x}=\mathbf{LayerNorm}(x+\mathbf{Drop}(\mathbf{MHA}(Q^{\prime},K^{\prime},V^{\prime})))$ (9)
@@
For ease of expression, we denote regular multi-head attention as $\mathbf{MHA}$ and multi-head attention with edge features as $\mathbf{MHA}_{edge}$. Our Scene-centric Feature Interaction can be denoted as follows:
$\hat{\mathcal{X}}^{(l)}=\mathbf{MHA}_{MA}(\hat{\mathcal{M}}^{(l-1)},\hat{\mathcal{X}}^{(l-1)})$ (10)
^$\hat{\mathcal{M}}^{(l)}=\mathbf{MHA}_{MM}(\hat{\mathcal{M}}^{(l-1)})$ (11)
^$\hat{\mathcal{M}}^{(l)}=\mathbf{MHA}_{AM}(\hat{\mathcal{X}}^{(l)},\hat{\mathcal{M}}^{(l-1)})$ (12)
$\hat{\mathcal{X}}^{(l)}=\mathbf{MHA}_{AA}(\hat{\mathcal{X}}^{(l-1)})$ (13)
@@
here, $l$ denotes the current layer index, and $\hat{\mathcal{X}}^{(0)}=\hat{\mathcal{X}}$ and ^$\hat{\mathcal{M}}^{(0)}=\hat{\mathcal{M}}$ are the initial inputs.
Our Agent-centric Feature Interaction can be denoted as follows:
$\hat{\mathcal{A}}^{(k)}=\mathbf{MHA}_{edge}(\hat{\mathcal{A}}^{(k-1)},\hat{\mathcal{L}}^{(k-1)},\hat{\mathcal{RPE}}^{(k-1)})$ (14)
@@
similarly we complete multiple rounds of interactions by looping through multiple levels of updates.
III-E Difficulty-guided decoder
To take advantage of the different prediction difficulties of different agents in the scenario, we developed the Difficulty Guided Feature Enhancement Module, which consists of a Difficulty Guided Trajectory Decoder and a Future Feature Interaction Module.

### v3

@@ Abstract
Trajectory prediction is crucial for autonomous driving, as it aims to forecast the future movements of traffic participants. Traditional methods usually perform holistic inference on the trajectories of agents, neglecting differences in prediction difficulty among agents. This paper proposes a novel Difficulty-Guided Feature Enhancement Network (DGFNet), which leverages the prediction difficulty differences among agents for multi-agent trajectory prediction. Firstly, we employ Spatio-temporal Feature Extraction to capture rich spatio-temporal features. Secondly, a Difficulty-Guided Decoder controls the flow of future trajectories into subsequent modules, obtaining reliable future trajectories. Then, feature interaction and fusion are performed through the Future Feature Interaction module. Finally, the fused actor features are fed into the Final Decoder to generate the predicted trajectory distributions for multiple participants. Experimental results demonstrate that our model achieves SOTA performance on the Argoverse 1&2 motion forecasting benchmarks. Ablation studies further validate the effectiveness of each module. Moreover, compared to the SOTA methods, our method balances trajectory prediction accuracy and real-time inference speed. The code is available at https://github.com/XinGP/DGFNet.
@@ I Introduction
Trajectory prediction is a crucial component of current autonomous driving systems. The objective of this task is to infer the future trajectory distribution of agents by considering their historical trajectories, dynamic and static scene information, and the interactions between these elements. By accounting for these complex factors, the future trajectory distribution can accurately reflect the agent’s motion trends, ensuring the proper functioning of the autonomous driving system.
@@
Previously, architectures of mainstream prediction models could be roughly divided into three steps: feature encoding, feature interaction, and feature decoding. Recently, some methods have improved the accuracy and robustness of prediction models through feature enhancement [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. For example, some studies model the future intentions of vehicles as latent variables and enhance features based on these intentions [1, 2, 3, 4, 10]. Other studies model intermediate latent variables in the form of trajectories, where the outputs of the initial decoders are one or more predicted trajectories [5, 6, 7]. In addition to network prediction methods, initial trajectories or intentions can also be obtained through planners or distance sampling [8, 9].
Although these methods significantly improve model performance compared to mainstream model architectures, they neglect the inherent heterogeneity in prediction difficulty among agents, i.e., the natural differences in prediction difficulty for different agents within the scene. In real driving scenarios, human drivers often predict the future behaviors of most agents subconsciously, relying on intuition [11]. However, when faced with agents that may pose risks and exhibit high prediction difficulty, human drivers tend to carefully consider their behavior and motivation, such as those with tracking loss or low-quality trajectories due to perception issues [12, 13, 14], or those exhibiting aggressive driving behaviors [15]. As illustrated in the Fig. 1, The primary difference in our pipeline is the addition of an intermediate module that generates reliable future trajectories, thus better prediction results can be achieved for agents with higher prediction difficulty.
Moreover, due to the transition in trajectory prediction tasks from single-agent to multi-agent scenarios, the research trend in scene representation has shifted from scene-centric to agent-centric approaches. However, embedding future trajectories & anchors into agent-centric scene representation methods is highly inconvenient. This is because agent-centric scene representation methods require pairwise-relative pose to describe spatial positional relationships between lanes, agents, and other elements [16, 17, 18, 19, 20]. Consequently, agent-centric encoded features cannot comprehend the coordinate information of future trajectories and align features accordingly. To address this, we aim to build a future feature enhancement network for multi-agent trajectory prediction by leveraging the strengths of both scene representation approaches.
Inspired by human drivers, we aim to investigate the effectiveness of prediction methods that follow the principle of easy-to-hard in multi-agent trajectory prediction tasks. The main contributions of this paper can be summarized as follows:
@@
The organization of this paper is as follows: Section II provides a detailed introduction to our pipeline; Section III analyzes the experimental details and results of our proposed model and compares it with several SOTA methods; finally, Section IV summarizes our research and discusses future related research directions.
II Method
II-A Overview
@@
(1)Spatio-temporal Feature Extraction. This component extracts spatio-temporal features separately from agent-centric historical trajectory information and scene-centric information using both agent-centric and scene-centric approaches. (2)Feature Interaction. Unlike Spatio-temporal Feature Extraction, this component is not compatible with both scenarios. However, feature interaction is not. The feature interaction component is categorized into two types: multi-head attention module and multi-head attention module with edge features. The two interaction methods are respectively applicable to scene-centric and agent-centric scenario representations. (3)Trajectory Decoder. This component is divided into a Difficulty-Guided Decoder and a Final Decoder. The Difficulty-Guided Decoder operates in the intermediate stage, while the Final Decoder generates the prediction results.
@@
Figure 3: Our Spatio-temporal Feature Extraction includes two sets of independent encoders, which extract features based on scene representations (bottom). Subsequently, the extracted features pass through their respective Feature Interaction modules to obtain interacted actor features. Finally, we obtain the predicted trajectories and their corresponding probabilities through the trajectory decoder with Future Feature Enhancement and Difficulty-Guided Decoder.
II-B Problem Formulation
Building on the majority of previous work, such as [21, 22, 1, 18, 15], we assume that upstream perceptual tasks can provide high-quality 2D trajectory tracking data for agents in a coordinate system. At the same time, localization and mapping tasks can provide accurate self-vehicle positioning data and comprehensive HD map data for the urban environment. In other words, for $j$ agents within a given scene, we obtain $x$ and $y$ positions, denoted as $\mathcal{X}^{0:j}_{-t_{h}:0}$, corresponding to the time stamp horizon $\left\{-t_{h},...,0,1,...,t_{f}\right\}$. The trajectory prediction task consists of predicting the future trajectories $\mathcal{Y}^{0:j}_{1:t_{f}}$ of the agents by using the HD map information(including road coordinates, topological links, traffic lights) within the given scene and the historical trajectories output by the tracking task. We represent the trajectory information of agents and lane information in vectorized form, where the agent trajectories include historical spatio-temporal sequences, and the map information includes lane coordinates, types, and traffic signal details.
@@ II-B Problem Formulation
II-C Spatio-temporal Feature Extraction
In the problem formulation, we explain the process of representing agent and high-definition (HD) map data as vectors, establishing a corresponding mapping between continuous trajectories, map annotations, and sets of vectors. This vectorization method allows us to encode trajectory and map information as vectors. Following the trajectory feature extraction method of LaneGCN [21], our trajectory feature extraction module mainly uses 1D CNN and downsampling techniques to encode the historical trajectories of all vehicles in the scene, thus obtaining encoded historical trajectory features. For the map tensor, we use the PointNet [23] to encode lane nodes and structural information to obtain map features. It is worth mentioning that, when processing scene-centric vector information, we keep vehicle historical trajectories and lane line information separate. Our view is that vehicle trajectories, as dynamic vectors, need to be distinguished from static maps to better capture the local motion characteristics of the vehicles. The extraction of historical trajectory features is described as follows:
$\begin{split}\mathcal{Z}_{i}&=\rm{Conv1d}(\rm{Res1d}(\mathcal{X}_{i})),\quad i=0,1,\dots,n_{\text{fpn}}-1,\\ \hat{\mathcal{X}}_{i}&=\rm{Inter}(\mathcal{Z}_{i+1},\rm{Scale}=2)+\mathcal{Z}_{i}\quad\end{split}$ (1)
after Res1d and Conv1d encoding, feature alignment is performed by scaling and interpolation operations. For the agent-centric trajectory tensor, the same feature extraction is used to obtain $\hat{\mathcal{A}}$. The process of map feature extraction is described as follows:
$\begin{split}{\mathcal{H}}_{1}&=\rm{PAB}_{1}\left(\rm{ReLU}(\rm{LayerNorm}(\mathcal{M}\textbf{W}_{p}+\textbf{b}_{p}))\right),\\ \hat{\mathcal{M}}&=\rm{PAB}_{2}({\mathcal{H}}_{1})\end{split}$ (2)
where $\mathcal{M}\in\mathbb{R}^{N_{\text{lane}}\times 10\times C_{\text{in}}}$ are the input lane features, $\textbf{W}_{p}\in\mathbb{R}^{C_{\text{in}}\times h}$ and $\textbf{b}_{p}\in\mathbb{R}^{h}$ are the weights and biases for the projection layer, and $\rm{PAB}$ represents the PointAggregateBlock. Similarly, we can obtain the agent-centric map feature $\hat{\mathcal{L}}$. For $\mathcal{RPE}$ we use $\rm{MLPs}$ for feature extraction to get $\hat{\mathcal{RPE}}$.
II-D Feature Interaction
Our intention was to utilize the global representation of scene-centric scenes to embed future trajectory features. Interestingly, we found that even without any future enhancement, simply concatenating the actor features after the two interactions provides a noticeable improvement. Specific results can be seen later in the ablation experiments. Here we will present the two interaction methods: multi-head attention module and multi-head attention module with edge features, respectively.
@@ II-D Feature Interaction
$\rm{MHA}(\textbf{Q},\textbf{K},\textbf{V})=\rm{Softmax}\left(\frac{\textbf{Q}\textbf{K}^{T}}{\sqrt{dim_{k}}}\right)\textbf{V},$
$\textbf{Q},\textbf{K},\textbf{V}=\textbf{W}^{q},\textbf{W}^{k},\textbf{W}^{v}$ (3)
where $\rm{MHA}$ denotes the Multi-Head Attention operation. Q, K, V are computed by linear projections $\textbf{W}^{q}$, $\textbf{W}^{k}$, $\textbf{W}^{v}$ applied to input vectors. The feature interaction module for carrying edge features is described as follows:
@@
$\begin{split}&\rm{MHA}(\textbf{Q}^{\prime},\textbf{K}^{\prime},\textbf{V}^{\prime})=\rm{Softmax}\left(\frac{\textbf{Q}^{\prime}\textbf{K}^{\prime T}}{\sqrt{dim_{k}}}\right)\textbf{V}^{\prime},\\ &\textbf{Q}^{\prime},\textbf{K}^{\prime},\textbf{V}^{\prime}=\textbf{Q}+\textbf{W}^{q}\text{edge}^{\prime},\textbf{K}+\textbf{W}^{k}\text{edge}^{\prime},\textbf{V}+\textbf{W}^{v}\text{edge}^{\prime}\end{split}$ (4)
@@
where $\textbf{Q}^{\prime},\textbf{K}^{\prime},\textbf{V}^{\prime}$ are the query, key, and value tensors after linear transformations that incorporate edge features, and $\text{edge}^{\prime}$ is the edge feature after linear transformation. Finally, the normalized output after attention and dropout application is described as follows:
$\hat{\textbf{x}}=\rm{LayerNorm}(\textbf{x}+\rm{Drop}(\rm{MHA}(\textbf{Q}^{\prime},\textbf{K}^{\prime},\textbf{V}^{\prime})))$ (5)
@@
For ease of expression, we denote regular multi-head attention as $\rm{MHA}$ and multi-head attention with edge features as $\rm{MHA}_{edge}$. Our Scene-centric Feature Interaction is described as follows:
$\begin{split}&\hat{\mathcal{X}}^{(l)}=\rm{MHA}_{MA}(\hat{\mathcal{M}}^{(l-1)},\hat{\mathcal{X}}^{(l-1)}),\\ &\hat{\mathcal{M}}^{(l)}=\rm{MHA}_{MM}(\hat{\mathcal{M}}^{(l-1)}),\\ &\hat{\mathcal{M}}^{(l)}=\rm{MHA}_{AM}(\hat{\mathcal{X}}^{(l)},\hat{\mathcal{M}}^{(l-1)}),\\ &\hat{\mathcal{X}}^{(l)}=\rm{MHA}_{AA}(\hat{\mathcal{X}}^{(l-1)})\end{split}$ (6)
@@
where $l$ denotes the current layer index, and $\hat{\mathcal{X}}^{(0)}=\hat{\mathcal{X}}$ and ^$\hat{\mathcal{M}}^{(0)}=\hat{\mathcal{M}}$ are the initial inputs. Our Agent-centric Feature Interaction is described as follows:
$\hat{\mathcal{A}}^{(k)}=\rm{MHA}_{edge}(\hat{\mathcal{A}}^{(k-1)},\hat{\mathcal{L}}^{(k-1)},\hat{\mathcal{RPE}}^{(k-1)})$ (7)
@@
similarly, we complete multiple rounds of interactions by looping through multiple levels of updates.
II-E Trajectory Decoder
As shown in Fig. 3, our Trajectory Decoder consists of three modules. First, the actor features from the scene-centric representation are passed into the Difficulty-Guided Decoder module to obtain reliable future trajectories and encode them. Then, through the Future Feature Enhancement module, the actor features from the agent-centric representation are expressed and aligned in the global coordinate system. Finally, the interacted features are passed into the Final Decoder to obtain the final prediction results.